# Advancing AI Research Assistants with Expert-Involved Learning

### v2

@@ 1 Introduction
The success of Foundation Models (FMs) (1), including Large Language Models (LLMs) and Large Multi-Modal Models (LMMs), has been demonstrated across various applications, garnering widespread public interest in research focused on these models (2, 3, 4, 5). A very important target of an FM is to understand various modalities and model them to enable multimodal reasoning (6, 7) as humans. These human skills give us the ability to understand what is being conveyed in a multimodal context, while at the same time being able to externalize it with our linguistic system for processing and refining the content of the text (8). In this work, we investigate how we can utilize FMs to help humans summarize long research papers as well as understand multimodal scientific context.
@@
There are various long documents in the biomedical domain, including examination reports, patient case data, research manuscripts, and others (9). These documents may also include figures as extra information to help us understand the corresponding content. However, due to the explosion of biomedical data nowadays, physicians and researchers are struggling more and more to understand and process such documents in biomedical areas (10). Also, the cost of hiring additional human resources to help process the data is high and therefore not optimal for the biomedical field (11, 12). At the same time, there has been evidence (13, 14, 15) to show that humans tend to engage in complicated reasoning processes and expand their knowledge when answering realistic questions. For example, it is common to use multiple examination results to determine the condition (16) during the diagnosis from a physician. Therefore, how to measure and optimize the reasoning capacity of FMs for complex multimodal data is an important and practical research problem (17). If an FM shows an excellent ability to summarize long texts as well as complicated figures with different panels, it will become a powerful research assistant in this field, which can shorten the time for physicians to make correct diagnoses and for researchers to understand the work of their peers. Previous research has investigated the ability of LLMs to summarize non-medical documents including news (18), codes (19), and reviewer comments (20, 21). Research has also been conducted to evaluate LLMs for summarizing short medical documents including clinical texts (12) and medical evidence (22). Meanwhile, the ability of LMMs to summarize scientific figures has also been explored (23, 24). These papers argue that some advanced models such as GPT-4 can outperform human experts in highlighting the abstract of certain documents or figures, while all models have their limitations due to either hallucinations or out-of-domain queries. Furthermore, most of the recent LLM or LMM evaluations have focused on comparisons among models but ignored the role of humans in evaluations. However, humans are very important references in evaluating the utilities of FMs for their purpose, and understanding the differences between FMs and humans can further ensure the reliability of these FMs for scientific research.
@@
As far as we know, comprehensive human-aware evaluations of LMMs for understanding long biomedical documents and biomedical figures remain unexplored. Although there are several LMM benchmarks on low-level knowledge quizzes, they cannot be applied to our problem due to the complexity of biomedical resources (25, 26, 27). In addition, due to the rapid development of LMMs, we expect more up-to-date evaluations and benchmarks in this time-sensitive domain (28). Most of the existing baselines also do not include human participators as an important metric. Researchers need to get fair and unbiased results carefully to avoid leakage of the test set. Furthermore, inspired by previous work focusing on generating scientific hypotheses from text-level information (29, 30), we are interested in understanding how multimodal biomedical data can help generate new research questions to accelerate scientific progress. Therefore, our LMM evaluations and improvement are challenging and meaningful, which can fill a gap in this advanced area.
@@ 2 Results
Overview of ARIEL To systemically assess the abilities of FMs in understanding research papers, we constructed two datasets: one from manuscripts and another from figures. Data were extracted from PubMed (31) and recently published top-tier journals, with no overlap to the cut-off dates of model training sets. We then developed a general pipeline for generating and automatically evaluating outputs from both closed- and open-source FMs on these datasets (Methods). For text summarization, a language model receives a manuscript and a prompt as input; for figure understanding, a multimodal model takes the figure, its description, and a prompt. In addition to automated metrics, we conducted human evaluation (Human-Eval) with 12 doctoral-level biomedical researchers who judged model outputs and their sources. We then analyzed all experimental results to identify the gaps between current FM capabilities and the requirements of an effective AI research assistant. Finally, we propose initial strategies to leverage FMs for summarizing documents, interpreting figures, and generating hypotheses. A overview of the ARIEL framework is shown in Figure 1.
@@
Datasets. All questions in our datasets are open-ended. For document summarization, we collected 2571 manuscripts via the PubMed API. Compared with existing benchmarks (32, 33, 34), ARIEL is both both domain-specific and more rigorous, reflecting the complexity of scientific language. We posed open-ended prompts rather than standard QA to increase difficulty. Manuscripts were split into training and test sets, with validation sets derived automatically (35). Topic analysis (Figure 1 (b)) shows dominance of COVID-19 and SARS-CoV-2, followed by healthcare and treatment, reflecting a focus on respiratory epidemics.
For figure understanding, we curated 10 high-quality figures from peer-reviewed publications (24, 25), paired with 88 open-ended questions spanning summarization, color identification, model design, rank identification, recommendation, and insight discovery (Figure 1 (c)). Compared with prior benchmarks (24, 25), our dataset prioritizes low noise and questions requiring reasoning. Human evaluation (Human-Eval) included all figures and a random sample of 10 manuscripts.
@@
Large-scale evaluations of long document summarization identify fine-tuned models as better performers. We benchmarked both open-source LLMs (ChatGLM4 (9B) (36), Gemma1 (8B) (37), Llama 3.1 (8B) (38), Mistral (7B) (39), Deepseek-v3 (40), GPT-OSS (20B) (41), GPT-OSS (120B) (41)) and closed-source LLMs (GPT-3.5 (42), GPT-4 (43), GPT-4o (44), Gemini-Flash-1.5 (45), and Claude-Sonnet-3.5 (46)). Performance was assessed with seven established metrics (BLEU, ROUGE-1/2/L, BERTScore, F1-Radgraph, MEDCON)(12, 47, 48, 49, 50, 51, 47). All of the metrics range from 0 to 100 and a higher value means a better output (Figure 2 (a), Methods). Figure 2 (b) shows the pipeline for LMM evaluation, which will be discussed later. To test whether task-specific adaptation improves performance, we fine-tuned the best-performing open-source model, ChatGLM4, using Low-Rank Adaptation (LoRA (52)).
Based on Figure 2 (c), we analyzed the distribution and variance of scores across all evaluation metrics. Overall, open-source models operating in zero-shot inference mode exhibited lower performance than closed-source models. Even advanced open-source systems, such as DeepSeek-V3 and the GPT-OSS series, achieved comparatively modest scores, while models like Gemma1, Llama 3.1, and Mistral often failed to generate meaningful outputs. These shortcomings are likely attributable to data constraints and limited context length. By contrast, ChatGLM4, Claude-Sonnet-3.5, and GPT-4 consistently ranked among the top three models for long-document summarization, aligning with prior studies on the use of LLMs as reviewers (53). Notably, ChatGLM4 surpassed other open-source models across all metrics and demonstrated competitive performance relative to closed-source systems. To further explore its capacity, we applied different prompting strategies and fine-tuned the model (ChatGLM_lora). As shown in Extended Data Figure 2, prompt engineering methods including meta prompting (54) and chain-of-thought reasoning (CoT) (6) had minimal impact on performance. However, fine-tuning substantially improved outcomes, with the adapted model outperforming all closed-source LLMs. This result suggests that domain-specific training can significantly enhance summarization quality and that fine-tuned, smaller-scale models can achieve strong task-specific performance. In Figure 2 (d), we display the average score of each metric for all the models, and the results of both performances and ranks are consistent with conclusions from Figure 2 (c).
@@
Interpretation of the human-generated results of long document summarization. We collected expert feedback on abstract generation for selected scientific papers, denoted as human-generated results. The participants were highly trained specialists in biology, and their evaluations provide critical insights into the divergence between human and LLM preferences when summarizing long documents. As described in the Dataset section, ten papers were sampled from the testing set, and seven participants (denoted H1–H7) contributed feedback. According to Figure 3 (a), the fine-tuned LLM, ChatGLM_lora, can still outperform the best human participator H4 significantly, supported by the p-value computed based on two-sided Wilcoxon rank-sum test (p-value=0.015). Moreover, we demonstrated the performance comparison stratified by different metrics, shown in Figure 3 (b). According to this figure, ChatGLM_lora outperformed other human participators in all of the metrics, which demonstrates that the improvement of fine-tuned LLM is comprehensive. Moreover, we compared the time cost and monetary cost brought by humans and models. Here the cost of machines is computed based on running inference with the fine-tuned ChatGLM model with a standard NVIDIA H100 GPU (per hour price: 0.37\$) (55). The cost of human participators is computed based on averaging time across different participators based on their reports, and the per-hour price is estimated based on 100,000\$ as annual salary (per hour price: 48.0\$) (56). According to Extended Data Figure 3 (a), using LLM can reduce both time cost and monetary cost. Overall, these results indicate that fine-tuned LLMs can generate abstracts more closely aligned with original texts than human experts, but fine-tuned LLMs also show an increase in output variability.
@@
Without test-time computation scaling, current LMMs fail to fully understand scientific figures in research papers. In this section, we assess the capacity of LMMs to interpret scientific figures. Human interpretation of such figures typically involves understanding both the overall context and the detailed content of individual panels, as well as deriving potential insights or biological discoveries. To evaluate these dimensions, we designed a set of questions targeting different levels of comprehension and included both open-source models (Qwen-VL-Chat (57), LLaVA-13B (58)) and closed-source models (Gemini-Flash-1.5 (45), GPT-4V, GPT-4o, Claude-Sonnet-3.5 (46), and o1 (59)). The o1 model, in particular, leverages test-time computation (60) to improve reasoning by employing search strategies to identify the most accurate reasoning path when generating responses.
@@
Finally, inspired by (61), we investigated whether LMMs are capable of producing reliable self-assessments of confidence. This analysis is critical for understanding the internal logic guiding model-generated content. Specifically, we asked o1, GPT-4V, and Claude-Sonnet-3.5 to assign confidence scores (ranging from 0 to 1, with higher values indicating greater confidence) to their answers across all questions. As shown in Extended Data Figure 6, neither GPT-4V nor Claude-Sonnet-3.5 produced meaningful confidence estimates, as their assigned scores were not significantly correlated with evaluation metrics (Spearman correlation, p = 0.11 and p = 0.14, respectively). In contrast, o1 exhibited a significant positive correlation between confidence and accuracy (p = 0.006), although inconsistencies remained, with the model assigning divergent confidence levels to responses with identical scores. Overall, these findings raise concerns about the reliability of confidence estimation in current LMMs and underscore the need for careful human oversight when employing them as assistants for scientific figure interpretation.
@@
Interpretation of the Human-Eval results of figure understanding. In this section, we compare the performance of human participants and LMMs on the figure understanding task, highlighting both similarities and differences in their outputs. To approximate the upper bound of LMM performance, we introduced two additional baselines for o1: the five-majority-vote variant (o1 (5 mv)) and the best-of-five variant (o1 (bo5)), the latter option serving as an oracle. We also had eight human experts (H1–H8) in the evaluation. According to Figure 5 (a), the performer with the best average accuracy in answering figure-related questions is still human, while LMMs with test-time computation with three different variations can only surpass two human participators. Moreover, major voting improves the accuracy of the LMM and reduces the difference in performance compared with the best human participators. This conclusion is supported by the p-value difference of the two-sided Wilcoxon rank-sum test. However, major voting consumes more tokens and therefore leads to higher costs. Therefore, we also explored how to utilize rationale engineering to improve the answer accuracy of LMM in the next section. In Figure 5 (b), we visualize the differences between the performances of LMMs and humans by categories. We found that o1 (5 mv) ranks higher than o1 in figure summarization and counting ability apparently, the LLMs might have more randomness in generating answers related to these tasks. Moreover, in all of the categories, the best performer is human, but the human participators also have in-group variation across different categories. However, according to Extended Data Figure 3 (b), LMMs still have advantages in reducing both the time cost and monetary cost for this task. Here the monetary cost of LMMs is computed based on the token price of API calling (62). These findings indicate that integrating human feedback remains valuable for improving LMM performance. However, to maximize benefits, careful selection of reference participants is necessary, prioritizing individuals with higher accuracy to ensure the reliability of training or evaluation data.
@@
We further conducted error analysis of generated answers, with error types defined following prior studies (63, 64). Figure 5 (d) illustrates a case in which all LLMs and most human participants produced incorrect responses. The task required describing the relationship between CPS values and the relative abundance of cell types across two regions. In this instance, o1 grouped the two regions together and emphasized shared patterns, whereas the correct human responses separated analyses by both region and cell type. This outcome suggests that LLMs often overlook fine-grained details and require improved capacity for figure-specific inference. Human errors in this example primarily stemmed from not stratifying the analysis by cell type, highlighting the intrinsic difficulty of extracting precise information from scientific figures. A contrasting example is shown in Figure 5 (e), where all LLMs failed but nearly all human participants answered correctly. The question asked participants to identify the method with the greatest variation across metrics. While LLMs incorrectly selected another option, their reasoning process appeared consistent with human logic, as both relied on interpreting boxplot widths. The discrepancy suggests that although LLMs followed the correct reasoning path, they failed in execution, whereas humans directly identified the correct method. All identified error categories are summarized in Supplementary File 1.
@@
Exploring the hypothesis-generation capacity of ARIEL framework with text-image paired inputs. Finally, we explore the potential of leveraging ARIEL to integrate multimodal understanding for scientific hypothesis generation, with the goal of inspiring novel directions in biomedical research. In this analysis, we reuse scientific images from the previous tasks but apply new prompts specifically designed for hypothesis generation. We investigate two modes: (1) a single-LLM mode using o1 as the base model, and (2) a Multi-LMM agent mode (HypoGenAgent), also based on o1. HypoGenAgent employs four distinct LMMs assigned to different roles, following a modified design from (29). Here the Analyst takes the multimodal inputs and perform quantitative analysis, the Engineer executes the experimental plan, the Scientist proposes new hypothesis, and the Criticist can verify the quality of generated hypothesis. An overview of the framework is provided in Figure 7 (a). To evaluate the quality of the generated hypothesis, we consider different aspects modified from (65) with a comprehensive cover for evaluating scientific insights:
@@
Each evaluation criterion was quantified on a scale from 0 to 100, where 0 indicates complete non-compliance and 100 indicates full compliance. To mitigate potential bias in the evaluation process, we employed Claude-Sonnet-3.7 as an automated judge (66), which scored the outputs from each generation mode (results shown in Figure 7 (b)). According to this figure, although HypoGenAgent (o1) has a higher score compared with o1 on average, we do not observe statistically significant differences in the hypothesis generated by two different modes (p-value>0.05>0.05 with two-sided Wilcoxon rank-sum test). To better understand the properties of the hypothesis generated by these modes, we select representative examples of generated hypotheses for in-depth analysis. Based on Figure 7 (c), we found that the hypothesis from o1 generally has a longer length and more figure-oriented information, while HypoGenAgent (o1) tends to generate hypotheses with shorter lengths and more general plans. Moreover, o1 tends to discuss more gene-function-level information while HypoGenAgent (o1)’s outputs contains therapeutic discovery. Therefore, there exists a completeness-conciseness balance when using ARIEL as a hypothesis generator, which may help researchers to design ideas that fit well with their requirements.

### v3

@@ 1 Introduction
The success of Foundation Models (FMs) (bommasani2021opportunities), including Large Language Models (LLMs) and Large Multi-Modal Models (LMMs), has been demonstrated across various applications, garnering widespread public interest in research focused on these models (singhal2023large, acosta2022multimodal, tu2024towards, han2025ateb). A very important target of an FM is to understand various modalities and model them to enable multimodal reasoning (wei2022chain, zhang2023multimodal) as humans. These human skills give us the ability to understand what is being conveyed in a multimodal context, while at the same time being able to externalize it with our linguistic system for processing and refining the content of the text (johnson2015logic). In this work, we investigate how we can utilize FMs to help humans summarize long research papers as well as understand multimodal scientific context.
@@
There are various long documents in the biomedical domain, including examination reports, patient case data, research manuscripts, and others (chatterjee2021informed). These documents may also include figures as extra information to help us understand the corresponding content. However, due to the explosion of biomedical data nowadays, physicians and researchers are struggling more and more to understand and process such documents in biomedical areas (wang2019big). Also, the cost of hiring additional human resources to help process the data is high and therefore not optimal for the biomedical field (wijeysundera2012techniques, van2024adapted). At the same time, there has been evidence (stacey2003complex, bystrom1995task, wilfredo2013acquiring) to show that humans tend to engage in complicated reasoning processes and expand their knowledge when answering realistic questions. For example, it is common to use multiple examination results to determine the condition (sanders2010every) during the diagnosis from a physician. Therefore, how to measure and optimize the reasoning capacity of FMs for complex multimodal data is an important and practical research problem (han2024hybridmind). If an FM shows an excellent ability to summarize long texts as well as complicated figures with different panels, it will become a powerful research assistant in this field, which can shorten the time for physicians to make correct diagnoses and for researchers to understand the work of their peers. Previous research has investigated the ability of LLMs to summarize non-medical documents including news (zhang2024benchmarking), codes (haldar2024analyzing), and reviewer comments (li2024chatcite, jin2024comprehensive). Research has also been conducted to evaluate LLMs for summarizing short medical documents including clinical texts (van2024adapted) and medical evidence (tang2023evaluating). Meanwhile, the ability of LMMs to summarize scientific figures has also been explored (wang2024scientific, roberts2024scifibench). These papers argue that some advanced models such as GPT-4 can outperform human experts in highlighting the abstract of certain documents or figures, while all models have their limitations due to either hallucinations or out-of-domain queries. Furthermore, most of the recent LLM or LMM evaluations have focused on comparisons among models but ignored the role of humans in evaluations. However, humans are very important references in evaluating the utilities of FMs for their purpose, and understanding the differences between FMs and humans can further ensure the reliability of these FMs for scientific research.
@@
As far as we know, comprehensive human-aware evaluations of LMMs for understanding long biomedical documents and biomedical figures remain unexplored. Although there are several LMM benchmarks on low-level knowledge quizzes, they cannot be applied to our problem due to the complexity of biomedical resources (xiao2024logicvista, yue2024mmmu, sun2024pathmmu). In addition, due to the rapid development of LMMs, we expect more up-to-date evaluations and benchmarks in this time-sensitive domain (white2024livebench). Most of the existing baselines also do not include human participators as an important metric. Researchers need to get fair and unbiased results carefully to avoid leakage of the test set. Furthermore, inspired by previous work focusing on generating scientific hypotheses from text-level information (qi2024large, gottweis2025towards), we are interested in understanding how multimodal biomedical data can help generate new research questions to accelerate scientific progress. Therefore, our LMM evaluations and improvement are challenging and meaningful, which can fill a gap in this advanced area.
@@ 2 Results
Overview of ARIEL To systemically assess the abilities of FMs in understanding research papers, we constructed two datasets: one from manuscripts and another from figures. Data were extracted from PubMed (white2020pubmed) and recently published top-tier journals, with no overlap to the cut-off dates of model training sets. We then developed a general pipeline for generating and automatically evaluating outputs from both closed- and open-source FMs on these datasets (Methods). For text summarization, a language model receives a manuscript and a prompt as input; for figure understanding, a multimodal model takes the figure, its description, and a prompt. In addition to automated metrics, we conducted human evaluation (Human-Eval) with 12 doctoral-level biomedical researchers who judged model outputs and their sources. We then analyzed all experimental results to identify the gaps between current FM capabilities and the requirements of an effective AI research assistant. Finally, we propose initial strategies to leverage FMs for summarizing documents, interpreting figures, and generating hypotheses. A overview of the ARIEL framework is shown in Figure 1.
@@
Datasets. All questions in our datasets are open-ended. For document summarization, we collected 2571 manuscripts via the PubMed API. Compared with existing benchmarks (chen2025benchmarking, wang2022squality, bai2023longbench), ARIEL is both both domain-specific and more rigorous, reflecting the complexity of scientific language. We posed open-ended prompts rather than standard QA to increase difficulty. Manuscripts were split into training and test sets, with validation sets derived automatically (zheng2024llamafactory). Topic analysis (Figure 1 (b)) shows dominance of COVID-19 and SARS-CoV-2, followed by healthcare and treatment, reflecting a focus on respiratory epidemics.
For figure understanding, we curated 10 high-quality figures from peer-reviewed publications (roberts2024scifibench, xiao2024logicvista), paired with 88 open-ended questions spanning summarization, color identification, model design, rank identification, recommendation, and insight discovery (Figure 1 (c)). Compared with prior benchmarks (roberts2024scifibench, xiao2024logicvista), our dataset prioritizes low noise and questions requiring reasoning. Human evaluation (Human-Eval) included all figures and a random sample of 10 manuscripts.
@@
Large-scale evaluations of long document summarization identify fine-tuned models as better performers. We benchmarked both open-source LLMs (ChatGLM4 (9B) (glm2024chatglm), Gemma1 (8B) (team2024gemma), Llama 3.1 (8B) (touvron2023llama), Mistral (7B) (jiang2023mistral), Deepseek-v3 (liu2024deepseek), GPT-OSS (20B) (gptoss2025), GPT-OSS (120B) (gptoss2025)) and closed-source LLMs (GPT-3.5 (papergpt3), GPT-4 (achiam2023gpt), GPT-4o (hurst2024gpt), Gemini-Flash-1.5 (team2023gemini), and Claude-Sonnet-3.5 (antrop2024)). Performance was assessed with seven established metrics (BLEU, ROUGE-1/2/L, BERTScore, F1-Radgraph, MEDCON)(van2024adapted, yim2023aci, papineni2002bleu, lin2004rouge, zhangbertscore, jain1radgraph, yim2023aci). All of the metrics range from 0 to 100 and a higher value means a better output (Figure 2 (a), Methods). Figure 2 (b) shows the pipeline for LMM evaluation, which will be discussed later. To test whether task-specific adaptation improves performance, we fine-tuned the best-performing open-source model, ChatGLM4, using Low-Rank Adaptation (LoRA (hulora)).
Based on Figure 2 (c), we analyzed the distribution and variance of scores across all evaluation metrics. Overall, open-source models operating in zero-shot inference mode exhibited lower performance than closed-source models. Even advanced open-source systems, such as DeepSeek-V3 and the GPT-OSS series, achieved comparatively modest scores, while models like Gemma1, Llama 3.1, and Mistral often failed to generate meaningful outputs. These shortcomings are likely attributable to data constraints and limited context length. By contrast, ChatGLM4, Claude-Sonnet-3.5, and GPT-4 consistently ranked among the top three models for long-document summarization, aligning with prior studies on the use of LLMs as reviewers (zhou-etal-2024-llm). Notably, ChatGLM4 surpassed other open-source models across all metrics and demonstrated competitive performance relative to closed-source systems. To further explore its capacity, we applied different prompting strategies and fine-tuned the model (ChatGLM_lora). As shown in Extended Data Figure 2, prompt engineering methods including meta prompting (suzgun2024meta) and chain-of-thought reasoning (CoT) (wei2022chain) had minimal impact on performance. However, fine-tuning substantially improved outcomes, with the adapted model outperforming all closed-source LLMs. This result suggests that domain-specific training can significantly enhance summarization quality and that fine-tuned, smaller-scale models can achieve strong task-specific performance. In Figure 2 (d), we display the average score of each metric for all the models, and the results of both performances and ranks are consistent with conclusions from Figure 2 (c).
@@
Interpretation of the human-generated results of long document summarization. We collected expert feedback on abstract generation for selected scientific papers, denoted as human-generated results. The participants were highly trained specialists in biology, and their evaluations provide critical insights into the divergence between human and LLM preferences when summarizing long documents. As described in the Dataset section, ten papers were sampled from the testing set, and seven participants (denoted H1–H7) contributed feedback. According to Figure 3 (a), the fine-tuned LLM, ChatGLM_lora, can still outperform the best human participator H4 significantly, supported by the p-value computed based on two-sided Wilcoxon rank-sum test (p-value=0.015). Moreover, we demonstrated the performance comparison stratified by different metrics, shown in Figure 3 (b). According to this figure, ChatGLM_lora outperformed other human participators in all of the metrics, which demonstrates that the improvement of fine-tuned LLM is comprehensive. Moreover, we compared the time cost and monetary cost brought by humans and models. Here the cost of machines is computed based on running inference with the fine-tuned ChatGLM model with a standard NVIDIA H100 GPU (per hour price: 0.37\$) (lambdagpucost). The cost of human participators is computed based on averaging time across different participators based on their reports, and the per-hour price is estimated based on 100,000\$ as annual salary (per hour price: 48.0\$) (incomecost). According to Extended Data Figure 3 (a), using LLM can reduce both time cost and monetary cost. Overall, these results indicate that fine-tuned LLMs can generate abstracts more closely aligned with original texts than human experts, but fine-tuned LLMs also show an increase in output variability.
@@
Without test-time computation scaling, current LMMs fail to fully understand scientific figures in research papers. In this section, we assess the capacity of LMMs to interpret scientific figures. Human interpretation of such figures typically involves understanding both the overall context and the detailed content of individual panels, as well as deriving potential insights or biological discoveries. To evaluate these dimensions, we designed a set of questions targeting different levels of comprehension and included both open-source models (Qwen-VL-Chat (bai2023qwen), LLaVA-13B (liu2023visual)) and closed-source models (Gemini-Flash-1.5 (team2023gemini), GPT-4V, GPT-4o, Claude-Sonnet-3.5 (antrop2024), and o1 (jaech2024openai)). The o1 model, in particular, leverages test-time computation (snell2024scaling) to improve reasoning by employing search strategies to identify the most accurate reasoning path when generating responses.
@@
Finally, inspired by (kadavath2022language), we investigated whether LMMs are capable of producing reliable self-assessments of confidence. This analysis is critical for understanding the internal logic guiding model-generated content. Specifically, we asked o1, GPT-4V, and Claude-Sonnet-3.5 to assign confidence scores (ranging from 0 to 1, with higher values indicating greater confidence) to their answers across all questions. As shown in Extended Data Figure 6, neither GPT-4V nor Claude-Sonnet-3.5 produced meaningful confidence estimates, as their assigned scores were not significantly correlated with evaluation metrics (Spearman correlation, p = 0.11 and p = 0.14, respectively). In contrast, o1 exhibited a significant positive correlation between confidence and accuracy (p = 0.006), although inconsistencies remained, with the model assigning divergent confidence levels to responses with identical scores. Overall, these findings raise concerns about the reliability of confidence estimation in current LMMs and underscore the need for careful human oversight when employing them as assistants for scientific figure interpretation.
@@
Interpretation of the Human-Eval results of figure understanding. In this section, we compare the performance of human participants and LMMs on the figure understanding task, highlighting both similarities and differences in their outputs. To approximate the upper bound of LMM performance, we introduced two additional baselines for o1: the five-majority-vote variant (o1 (5 mv)) and the best-of-five variant (o1 (bo5)), the latter option serving as an oracle. We also had eight human experts (H1–H8) in the evaluation. According to Figure 5 (a), the performer with the best average accuracy in answering figure-related questions is still human, while LMMs with test-time computation with three different variations can only surpass two human participators. Moreover, major voting improves the accuracy of the LMM and reduces the difference in performance compared with the best human participators. This conclusion is supported by the p-value difference of the two-sided Wilcoxon rank-sum test. However, major voting consumes more tokens and therefore leads to higher costs. Therefore, we also explored how to utilize rationale engineering to improve the answer accuracy of LMM in the next section. In Figure 5 (b), we visualize the differences between the performances of LMMs and humans by categories. We found that o1 (5 mv) ranks higher than o1 in figure summarization and counting ability apparently, the LLMs might have more randomness in generating answers related to these tasks. Moreover, in all of the categories, the best performer is human, but the human participators also have in-group variation across different categories. However, according to Extended Data Figure 3 (b), LMMs still have advantages in reducing both the time cost and monetary cost for this task. Here the monetary cost of LMMs is computed based on the token price of API calling (openaiapitoken). These findings indicate that integrating human feedback remains valuable for improving LMM performance. However, to maximize benefits, careful selection of reference participants is necessary, prioritizing individuals with higher accuracy to ensure the reliability of training or evaluation data.
@@
We further conducted error analysis of generated answers, with error types defined following prior studies (hua2025inductionbench, miner2024scheherazade). Figure 5 (d) illustrates a case in which all LLMs and most human participants produced incorrect responses. The task required describing the relationship between CPS values and the relative abundance of cell types across two regions. In this instance, o1 grouped the two regions together and emphasized shared patterns, whereas the correct human responses separated analyses by both region and cell type. This outcome suggests that LLMs often overlook fine-grained details and require improved capacity for figure-specific inference. Human errors in this example primarily stemmed from not stratifying the analysis by cell type, highlighting the intrinsic difficulty of extracting precise information from scientific figures. A contrasting example is shown in Figure 5 (e), where all LLMs failed but nearly all human participants answered correctly. The question asked participants to identify the method with the greatest variation across metrics. While LLMs incorrectly selected another option, their reasoning process appeared consistent with human logic, as both relied on interpreting boxplot widths. The discrepancy suggests that although LLMs followed the correct reasoning path, they failed in execution, whereas humans directly identified the correct method. All identified error categories are summarized in Supplementary File 1.
@@
Exploring the hypothesis-generation capacity of ARIEL framework with text-image paired inputs. Finally, we explore the potential of leveraging ARIEL to integrate multimodal understanding for scientific hypothesis generation, with the goal of inspiring novel directions in biomedical research. In this analysis, we reuse scientific images from the previous tasks but apply new prompts specifically designed for hypothesis generation. We investigate two modes: (1) a single-LLM mode using o1 as the base model, and (2) a Multi-LMM agent mode (HypoGenAgent), also based on o1. HypoGenAgent employs four distinct LMMs assigned to different roles, following a modified design from (qi2024large). Here the Analyst takes the multimodal inputs and perform quantitative analysis, the Engineer executes the experimental plan, the Scientist proposes new hypothesis, and the Criticist can verify the quality of generated hypothesis. An overview of the framework is provided in Figure 7 (a). To evaluate the quality of the generated hypothesis, we consider different aspects modified from (zhong2023goal) with a comprehensive cover for evaluating scientific insights:
@@
Each evaluation criterion was quantified on a scale from 0 to 100, where 0 indicates complete non-compliance and 100 indicates full compliance. To mitigate potential bias in the evaluation process, we employed Claude-Sonnet-3.7 as an automated judge (chen2024mllm), which scored the outputs from each generation mode (results shown in Figure 7 (b)). According to this figure, although HypoGenAgent (o1) has a higher score compared with o1 on average, we do not observe statistically significant differences in the hypothesis generated by two different modes (p-value>0.05>0.05 with two-sided Wilcoxon rank-sum test). To better understand the properties of the hypothesis generated by these modes, we select representative examples of generated hypotheses for in-depth analysis. Based on Figure 7 (c), we found that the hypothesis from o1 generally has a longer length and more figure-oriented information, while HypoGenAgent (o1) tends to generate hypotheses with shorter lengths and more general plans. Moreover, o1 tends to discuss more gene-function-level information while HypoGenAgent (o1)’s outputs contains therapeutic discovery. Therefore, there exists a completeness-conciseness balance when using ARIEL as a hypothesis generator, which may help researchers to design ideas that fit well with their requirements.